Solve the separable differential equation for.



dy/dx= [1+x] divided by [xy^15]






Use the following initial condition: y(1)=5


y^16=?



.


1
Expert's answer
2022-03-10T18:18:16-0500

Let us solve the separable differential equation

"\frac{dy}{dx}=\frac {1+x}{xy^{15}}"

which is equivalent to

"y^{15}dy=\frac {1+x}{x}dx"

It follows that

"\int y^{15}dy=\int\frac {1+x}{x}dx=\int(\frac{1}x+1)dx."

Therefore, "\frac{y^{16}}{16}=\ln|x|+x+C."

Since "y(1)=5," we get that "\frac{5^{16}}{16}=1+C," ànd hence "C=\frac{5^{16}}{16}-1."

We conclude that the solution is

"\frac{y^{16}}{16}=\ln|x|+x+\frac{5^{16}}{16}-1" or

"y^{16}=16\ln|x|+16x+5^{16}-16"